4^61+4^62+4^63+4^64 is divisible by [a]3[b]10[c]11[d]13

Grade:12

5 Answers

Shilpi Suman
13 years ago

Hi !!

Please define a,b,c and d, in the above question.

Raj
7 years ago
(461 + 462 + 463 + 464) = 461 x (1 + 4 + 42 + 43) = 461 x 85. = 460 x (4 x 85). = (460 x 340), which is divisible by 10. 
Avhinash Yadav
6 years ago
461+462+463+464=4(61+62+63+64)=4(250) because of divisibility rule  this no is divisible by 10 .so the answer is 10
Amaljith
4 years ago
4^odd= end with 4
4^even= end with 6
So 4^61+4^62+4^63+4^64= xxxxx4+xxxxx6+xxxx4+xxxxx6
=xxxxxxxx0
End with zero
So divisible by 10
Ashutosh bhai ji
one year ago
(4^61+4^62+4^63+4^64) 
=1+2+3+4=10
Sabhi ghaton ki ikai ke ank jodne per Jo Yog Aata Hai vah 10 se vibhajit hai to Puri sankhya bhi 10 se vibhajit Hogi